Noah and Ben are running in the Walker Street Fun Run. Noah runs at 7 mph, Ben at 5 mph. If they start at the same time, how long will it be before they are 1/3 mile apart?

Answer by stanbon(75887)

Noah and Ben are running in the Walker Street Fun Run. Noah runs at 7 mph, Ben at 5 mph. If they start at the same time, how long will it be before they are 1/3 mile apart?
------------------------
Noah DATA:
Rate = 7 mph ; time = x hrs. ; distance = 7x miles
-------------------------
Ben DATA:
Rate = 5 mph ; time = x hrs. ; distance = 5x miles
----------------
EQUATION:
Noah distance - Ben distance = (1/3) mile
7x - 5x = (1/3)
2x = 1/3
x = 1/6 hr = 10 minutes
====================
Cheers,
Stan H.